OpenStudy (anonymous):

tan^2Θ = ((1-cos(2Θ))/((1+cos(2Θ)) PROVE THE IDENTITY!

OpenStudy (anonymous):



OpenStudy (loser66):

tan^2 = sin^2/cos^2 = right hand side Dat sit

OpenStudy (anonymous):

wait HOW is the right side sin^2Θ/cos^2Θ

OpenStudy (anonymous):

????

OpenStudy (loser66):

how your nick is music? tan= sin/cos so , tan^2 = sin^2/cos^2 I don't know how to answer your How :)

OpenStudy (anonymous):

um no how your explanation is right.

OpenStudy (anonymous):

how is that equalto the right side

OpenStudy (loser66):

cos (2x) = 1-2sin^2 x--> sin^2x = (1-cos(2x))/2 that's numerator cos(2x) = 2cos^2-1 do the same to get denominator

OpenStudy (anonymous):

yeh. then wut

OpenStudy (anonymous):

i get it till there. then what do you do

OpenStudy (anonymous):

can you just show me all the steps so its clear, pls?

OpenStudy (loser66):

the proof is done. nothing to say. I am not on the computer. It 's hard to type much moe

OpenStudy (anonymous):

I'll gladly help. First off, is this the right side of the original equation?: \[\frac{ 1-\cos^2x }{ 1+\cos^2x }\]

OpenStudy (anonymous):

the right side is: 1- cos (2Θ) / (1+cos(2Θ) its cos 2theta

OpenStudy (anonymous):

Sorry for the late response. The double-angle formula states that \[\cos2\theta=\cos^2\theta-\sin^2\theta\]Therefore, the right side can be simplified to \[\frac{ 1-(\cos^2\theta-\sin^2\theta) }{ 1+(\cos^2\theta-\sin^2\theta) }\]And then further to \[\frac{ 1-\cos^2\theta+\sin^2\theta }{ 1+\cos^2\theta-\sin^2\theta }\]Using the pythagorean identity (sin^2x + cos^2x = 1, thus sin^2x = 1 - cos^2x and so on), it can be simplified again to \[\frac{ \sin^2\theta+\sin^2\theta }{ \cos^2\theta+\cos^2\theta }\]Now, we can do simple addition. \[\frac{ 2\sin^2\theta }{ 2\cos^2\theta }\]The 2 in the numerator and the 2 in the denominator can cancel. The ratio identity (also known as reciprocal identity) states that tanx is equal to sinx / cosx. This rule also applies to tan^2, which is equal to sin^2x / cos^2x. We can use this rule to take the expression out of fraction form. \[\tan^2\theta\]Now both sides are equal!